# Improving thermal state preparation of Sachdev-Ye-Kitaev model with reinforcement learning on quantum hardware

### v1

@@ Abstract
The Sachdev-Ye-Kitaev (SYK) model, known for its strong quantum correlations and chaotic behavior, serves as a key platform for quantum gravity studies. However, variationally preparing thermal states on near-term quantum processors for large systems ($N>12$, where $N$ is the number of Majorana fermions) presents a significant challenge due to the rapid growth in the complexity of parameterized quantum circuits. This paper addresses this challenge by integrating reinforcement learning (RL) with convolutional neural networks, employing an iterative approach to optimize the quantum circuit and its parameters. The refinement process is guided by a composite reward signal derived from entropy and the expectation values of the SYK Hamiltonian. This approach reduces the number of CNOT gates by two orders of magnitude for systems $N\geq 12$ compared to traditional methods like first-order Trotterization. We demonstrate the effectiveness of the RL framework in both noiseless and noisy quantum hardware environments, maintaining high accuracy in thermal state preparation. This work contributes to the advancement of a scalable, RL-based framework with applications for computations of thermal out-of-time-order correlators in quantum many-body systems and quantum gravity studies on near-term quantum hardware.
@@ 1 Introduction
The holographic duality [1] establishes a profound connection between a specific class of quantum field theories in $d$ dimensions and quantum gravity in $d+1$ dimensions. This strong/weak duality allows the exploration of strongly coupled field theories via classical supergravity and vice versa. Despite its power, there are no known cases where both sides of the duality can be simultaneously and analytically solved. The Sachdev-Ye-Kitaev (SYK) model, a simplified variant of the Sachdev-Ye model [2], was introduced in a series of seminal talks by Kitaev [3, 4]. This model exhibits holographic properties and is particularly amenable to study in the strong-coupling limit.
The SYK model consists of $N$ Majorana fermions in $0+1$ dimensions with random couplings between $q$ fermions, drawn from a Gaussian distribution with zero mean and variance proportional to $J^{2}/N^{q-1}$. In the large-$N$ and low-temperature limit $N\gg\beta J\gg 1$, where $\beta$ is the inverse temperature and $J$ the disorder strength—the model displays an approximate conformal symmetry and is related to near-extremal black holes with a nAdS2 (near-AdS2) geometry. Meanwhile, the SYK model saturates the chaos bound [5], a key feature of holographic systems. Moreover, SYK model is a prominent example of a fast scrambler [6], known for its logarithmic growth of complexity and its saturation of the bounds proposed by the Brown-Susskind conjecture regarding quantum information scrambling in black holes [7]. Its computational tractability in $0+1$ dimensions has enabled extensive numerical studies, with systems comprising up to 60 Majorana fermions analyzed [8, 9]. Over the past decade, the SYK model has been extensively studied in both condensed matter and high-energy physics [10, 11, 12]. For a comprehensive review, see refs. [13, 14, 15].
Given its significance in holography, detailed investigations of both pure and thermal (Gibbs) states of the SYK model are imperative. The ground state exhibits volume-law entanglement entropy with a known coefficient [16], rendering it classically challenging to prepare [17]. Additionally, a defining feature of the SYK model is the behavior of thermal four-point correlators, particularly out-of-time-order correlators (OTOCs), which saturate the Lyapunov exponent in the large-$N$ and low-temperature regime [5]. Accurate computation of these OTOCs requires preparing the corresponding thermal state, which is a central focus of this study. While low-energy states have been explored via large-scale classical simulations, our research aims to advance the state-of-the-art by employing hybrid quantum-classical algorithms [18, 19].
@@
Recent efforts have applied various quantum simulation techniques to the SYK model, including digital simulation [20], teleportation protocols [21], computation of bosonic correlation functions [22], and ground-state preparation of coupled SYK models [23, 24]. Notably, ref. [25] employed a variational quantum algorithm to prepare thermal states of dense and sparse SYK models for $6\leq N\leq 12$ Majorana fermions across a range of temperatures. Which is an extension of ref. [26], the authors focus on the real-time dynamics of the SYK model using quantum hardware. However, this approach faces significant challenges, including the rapid growth in the number of parameters and controlled-NOT (CNOT) gates with increasing $N$. These limitations restricted simulations to $N\leq 12$, as larger systems incurred prohibitive computational overhead.
Figure 2: The CNOT gates in the PQC2 are $10^{2}$-fold improved beyond $N=12$ if the thermal state is prepared in the reinforcement learning framework compared to the real-time dynamics obtained by a first-order Trotterization [26]. The resource count is obtained in a noiseless scenario where both the environment and the agent are noiseless. Using the reward signal presented in Eq. 4 (where the signal constitutes the error in free energy, hence “Free energy reward”) and in Eq. 5 (where the signal constitutes the error in free energy and the thermal state fidelity, hence the “Free energy and fidelity reward”) and for different temperatures ($1/\beta$). The “ours” denotes the variational thermal state preparation under the reinforcement learning framework in Fig. 1.
In this work, we address these challenges by integrating reinforcement learning (RL) with convolutional neural networks (CNNs) [27] into the variational quantum algorithm (VQA) framework. RL has recently emerged as a powerful tool for addressing quantum computing problems on near-term quantum devices [28, 29, 30, 31, 32, 33, 34]. In this approach, quantum circuits are modeled as sequences of actions governed by a trainable policy. The VQA cost function, independently optimized using a classical optimizer, provides periodic feedback that contributes to the final reward signal. This reward signal updates the policy, maximizing expected returns and guiding the selection of optimal actions for subsequent steps.
@@
The remainder of the sections is organized as follows: in Sec. 2 we elaborate on the preliminaries, such as the problem statement and the variational thermal state preparation algorithm. This is followed by an introduction to the variational preparation of the thermal states in the RL framework in Sec. 3. In Sec. 4 we thoroughly discuss the main results of the paper, including the efficiency of the RL framework in noiseless and realistic noisy scenarios. Finally in Sec. 5 we conclude the paper and discuss possible future work following the success of this paper.
@@ 2 Preliminaries
2.1 The problem: Dense SYK model
The Hamiltonian for the full or dense SYK model with $N$ Majorana fermions and $q$-fermion interaction terms is:
$H=\frac{(i)^{q/2}}{q!}\sum_{i,j,k,\cdots,q=1}^{N}J_{i_{1}i_{2}\ldots i_{q}}~{}~{}\chi_{i}\chi_{j}\chi_{k}\cdots\chi_{q},$ (1)
where $\chi_{i}$ are Majorana fermions satisfying the anticommutation relation $\{\chi_{i},\chi_{j}\}=\chi_{i}\chi_{j}+\chi_{j}\chi_{i}=\delta_{ij}$. We focus on the case where $q=4$, which corresponds to random all-to-all quartic interactions averaged over disorder. The random couplings $J_{i_{1}i_{2}i_{3}i_{4}}$ are real-valued and are sampled from a Gaussian distribution with the following properties for the mean and variance:
$\overline{J_{ijkl}}=0,\;\;\text{and}\;\;\overline{J_{ijkl}^{2}}=\frac{3!J^{2}}{N^{3}},$ (2)
respectively. We set $J=1$ in this work. The dimension of the Hilbert space is $\text{dim}(\mathcal{H})=2^{N/2}$ where $n=N/2$ is the number of qubits is, we have $N=2n$ since two Majorana fermions can be represented by one complex spinless fermion, which can be represented by a single qubit. The model can also be considered for $q>4$ and is solvable in the large $q$ limit [10]. However, here, we will restrict ourselves to the version given in Eq. (1) with $q=4$. For this work, we choose dimensionless parameters and refer to $\beta J$ simply as $\beta$.
One of the challenges of the dense SYK model is that the number of terms in the Hamiltonian grows as $\sim N^{4}$. To reduce the computational cost of classical or quantum simulations of the SYK model, a sparsified version was introduced in Ref. [37, 38].
Due to the computational overhead of the dense SYK model and the rapidly increasing number of CNOT gates in the PQC2 with increasing $N$, the authors in ref. [25] could not extend the task of thermal state preparation of the SYK model beyond 12 Majorana fermions using the variational quantum algorithm framework. We tackle this computational overhead by employing a reinforcement learning (RL) framework with an $\epsilon$-greedy policy. Meanwhile, the RL agent is defined with a 3D convolutional neural network that takes a 3D tensor, encodes the quantum circuit as input.
@@ 2.2 Variational thermal state preparation
Thermal states, represented by a density matrix $\rho_{\beta}$, are critical for studying quantum systems at finite temperatures. These states are challenging to prepare compared to ground states ($\beta\to\infty$) or high-temperature states ($\beta\to 0$), which are computationally simpler. The finite inverse temperature regime ($\beta$) is particularly important for analyzing quantum many-body systems such as the SYK model.
To prepare a thermal state, we minimize the Helmholtz free energy $F$, given by
$F=\langle H\rangle_{\beta}-TS_{v},$ (3)
where $T=1/\beta$ is the temperature, $\langle H\rangle_{\beta}=\text{Tr}[\rho_{\beta}H]$ is the energy expectation value, where $\rho_{\beta}$ is the thermal state at temperature $T$, and ${S}_{v}=-\text{Tr}[\rho_{\beta}\log\rho_{\beta}]$ is the von Neumann entropy. We adopt the algorithm from Ref. [39] whose subroutines are defined in the following.
@@
  • Measurement and entropy computation: Apply the unitary transformation PQC1($\vec{\theta}$), where ${\vec{\theta}}$ are the variational parameters. The resulting quantum state is |$\ket{\psi({\vec{\theta}})}=\text{PQC}_{1}({\vec{\theta}})\ket{0}^{\otimes n}$. Then measure qubits in the computational basis and compute the Shannon entropy:
    $S=-\sum_{i=1}^{2^{n}}p_{i}({\vec{\theta}})\log p_{i}({\vec{\theta}}),$
    where $p_{i}({\vec{\theta}})=|a_{i}({\vec{\theta}})|^{2}$ represents the probability of the $i^{\mathrm{th}}$ basis state.
@@
  • Initialization: Use the measurement outcome from PQC1 as the initial state for PQC2.
  • Ansatz construction: Employ a problem-specific or quantum hardware-inspired ansatz [40] with gates either derived from the trotterized time evolution operator or obtained from the native gateset from quantum hardware.
@@
    $\rho_{\mathrm{PQC}_{2}}=\sum_{i}|a_{i}({\vec{\theta}})|^{2}\ket{\Psi_{i}({\vec{\phi}})}\bra{\Psi_{i}({\vec{\phi}})},$
    where |$\ket{\Psi_{i}({\vec{\phi}})}=\text{PQC}_{2}({\vec{\phi}})\ket{b_{i}}$.
@@ 3 Variational thermal state preparation with reinforcement learning
We give an overview of the variational thermal state preparation of the SYK model within the reinforcement learning (RL) framework, for details of the framework see Appendix A. The algorithm is illustrated in Fig. 1. The framework is used to estimate the parameters of the PQC1 and the structure as well as parameters of the PQC2 elaborated in Sec. 2.2. Wherein we present state, the action representations, and the reward function used in this work.
@@
In the hybrid quantum-classical environment, the agent begins each episode with PQC1 (which includes a layer of RZ-RY-RZ rotations on each qubit followed by CNOT entangling gate), then it sequentially adds gates to the PQC2 (which starts as an empty circuit at the beginning of each episode) until the maximum number of actions is reached. At each step, the state is represented by the current circuit, i.e., PQC1 combined with PQC2, and the action is to append a gate. To enable deep RL methods, states and actions are encoded in a format suitable for neural networks. Specifically, each state is represented by a tensor-based binary encoding of the circuit’s gate structure, as described in Appendix B. The action space consists of CNOT gates and 1-qubit parameterized rotation gates (RX, RY, and RZ). Notably, we omit the continuous parameters of the rotation gates, focusing instead on the estimated energy of the circuit to define the state representation for the RL agent.
@@
The reinforcement learning (RL) agent utilizes a 3D convolutional neural network (CNN), as the quantum circuit is encoded as a 3D tensor (refer to Appendix B). In prior studies [29, 41, 28, 42], this 3D representation was typically flattened into a 1D format to be processed by a fully connected feedforward neural network (FNN) [35, 36]. However, flattening 3D data into 1D can result in the loss of critical spatial information, diminished model expressiveness, and potentially suboptimal performance. Architectures like 3D CNNs that can directly process the 3D structure are often preferable when working with volumetric data. If not flattened when operating on a structure like a 3D-CNN [43, 44], the neural net receives a dataset with distinguishable spatial information. In the case of the 3D binary encoding, where the X-axis represents the position of gates in PQC2, the Y-axis denotes the type of gate, and the Z-direction is the depth of PQC2. In Appendix. C we elaborately show that directly feeding higher-dimensional data through a 3D-CNN provides improved training, leading to lower approximation error with more compact parameterized quantum circuits compared to processing flattened data through an FNN. The hypermeters settings of the agent is discussed in Appendix F
@@
$r=\begin{cases}5,&\text{if }F_{t}(\beta=\text{const.},\vec{\theta},\vec{\phi})<\zeta_{F},\\ -5,&\text{if }F_{t}(\beta=\text{const.},\vec{\theta},\vec{\phi})<\zeta_{F}\;\text{and,}\;\text{step no.}=D_{\text{max}},\\ E_{\text{term}},&\text{otherwise}.\end{cases}$ (4)
@@
In Sec. 4 we show that using the reward Eq. 4 we can achieve a very good approximation of to the free energy, but it cannot provide as good an approximation to the energy expectation of the SYK Hamiltonian and the entropy. Hence we engineer the reward signal to incorporate the accuracy of the prepared thermal state by calculating its fidelity from the target thermal state after PQC2. Hence the modified reward function is given by
$r=\begin{cases}5,&\text{if }F_{t}(\beta=\text{const.},\vec{\theta},\vec{\phi})<\zeta_{F}\;\text{and,}\;\textrm{Fid}(\rho({\vec{\phi}}))\geq\zeta_{\textrm{Fid.}}\\ -5,&\text{if }\textrm{Fid}(\rho({\vec{\phi}}))<\zeta_{\textrm{Fid}}\;\text{and,}\;\text{step no.}=D_{\text{max}},\\ 0.6\times E_{\text{term}}+0.4\times\textrm{Fid}_{\text{term}},&\text{otherwise},\end{cases}$ (5)
Where $\textrm{Fid}_{\text{term}}=2\times\textrm{Fid}(\rho({\vec{\phi}}))-1$, $\textrm{Fid}(\rho({\vec{\phi}}))={\mbox{Tr}}\Big{(}\sqrt{\rho_{\beta}^{1/2}\rho(\vec{\phi})~{}\rho_{\beta}^{1/2}}\Big{)}$ and $\rho_{\beta}$ represents the density matrix that corresponds to the true thermal state. $\zeta_{\textrm{Fid}}$ is the user-defined threshold on fidelity, throughout the paper we choose $\zeta_{\textrm{Fid}}=0.9$.

### v2

@@ Abstract
The Sachdev-Ye-Kitaev (SYK) model, known for its strong quantum correlations and chaotic behavior, serves as a key platform for quantum gravity studies. However, variationally preparing thermal states on near-term quantum processors for large systems ($N>12$, where $N$ is the number of Majorana fermions) presents a significant challenge due to the rapid growth in the complexity of parameterized quantum circuits. This paper addresses this challenge by integrating reinforcement learning (RL) with convolutional neural networks, employing an iterative approach to optimize the quantum circuit and its parameters. The refinement process is guided by a composite reward signal derived from entropy and the expectation values of the SYK Hamiltonian. This approach reduces the number of CNOT gates by two orders of magnitude for systems $N\geq 12$ compared to traditional methods like first-order Trotterization. We demonstrate the effectiveness of the RL framework in both noiseless and noisy quantum hardware environments, maintaining high accuracy in thermal state preparation. This work advances a scalable, RL-based framework with applications for quantum gravity studies and out-of-time-ordered thermal correlators computation in quantum many-body systems on near-term quantum hardware. The code is available at solving_SYK_with_RL repository.
@@ 1 Introduction
The holographic duality [1] establishes a profound connection between a specific class of quantum field theories in $d$ and quantum gravity in $d+1$ dimensions. This strong/weak duality allows the exploration of strongly coupled field theories via classical supergravity and vice versa. Despite its power, there are no known cases where both sides of the duality can be simultaneously and analytically solved. The Sachdev-Ye-Kitaev (SYK) model, a simplified variant of the Sachdev-Ye model [2], was introduced in a series of seminal talks by Kitaev [3, 4]. This model exhibits holographic properties and is particularly amenable to study in the strong-coupling limit.
The SYK model describes a system of $N$ Majorana fermions in a quantum mechanical setting (i.e., in $0+1$ dimensions), where the interaction involves random couplings among $q$ distinct fermions. These random couplings, denoted by $J_{i_{1}i_{2}\ldots i_{q}}$, are independent Gaussian variables with zero mean and variance given by
$\langle J_{i_{1}i_{2}\ldots i_{q}}^{2}\rangle=\frac{J^{2}(q-1)!}{N^{q-1}}.$
In the regime where $N$ is very large and the temperature is low ($N\gg\beta J\gg 1$, with $\beta$ the inverse temperature and $J$ the characteristic interaction strength), the model displays an approximate conformal symmetry and is related to near-extremal black holes with an nAdS2 (near-AdS2) geometry. Meanwhile, the SYK model saturates the chaos bound [5], a key feature of holographic systems. Moreover, SYK model is a prominent example of a fast scrambler [6], known for its logarithmic growth of complexity and its saturation of the bounds proposed by the Brown-Susskind conjecture regarding quantum information scrambling in black holes [7]. Its computational tractability in $0+1$ dimensions has enabled extensive numerical studies, with systems comprising up to 60 Majorana fermions analyzed [8, 9]. Over the past decade, the SYK model has been extensively studied in both condensed matter and high-energy physics [10, 11, 12]. For a comprehensive review, see refs. [13, 14, 15].
Owing to its pivotal role in holographic duality, rigorous analysis of both vacuum and finite-temperature states in the SYK model becomes essential. The ground state displays extensive entanglement entropy scaling linearly with system size, with coefficients explicitly computed in Ref. [16] – a property making it classically intractable to simulate [17]. A hallmark characteristic emerges in thermal correlation functions: out-of-time-order correlators (OTOCs) exhibit maximal quantum Lyapunov exponents $\lambda_{L}=2\pi/\beta$ at low temperatures ($\beta J\to\infty$) in the large-$N$ limit [5]. This chaotic behavior necessitates precise preparation of thermal states, a key objective of this work. While prior studies have probed low-energy physics through classical simulation, we propose leveraging variational quantum algorithms [18, 19] to overcome these limitations and enable quantum simulations of nonperturbative dynamics.
@@
While condensed matter systems, such as topological interfaces [20]-naturally realize large-$N$ SYK-like physics, and classical approaches like exact diagonalization can currently handle up $N=34$ to 0 + 1-dimensional models and have been studied up to $N=60$ Majorana fermions [9], we see quantum simulation as a promising alternative route. Instead of trying to outdo classical computation, our focus is on harnessing the unique strengths of quantum hardware, like precise control over parameters and direct access to non-equilibrium dynamics [21, 22]. By developing efficient state preparation techniques tailored for near-term devices, we aim to explore regimes that are out of reach for ED, such as open-system dynamics or much larger $N$. In this way, our approach complements classical methods and sets the stage for tackling high-$N$ problems that could reveal new physics as quantum technology continues to advance.
Recent efforts have applied various quantum simulation techniques to the SYK model, including digital simulation [23], teleportation protocols [24], bosonic correlation computation [22], and the preparation of the ground state of the coupled SYK model [25, 26]. Notably, ref. [27] employed a variational quantum algorithm to prepare thermal states of dense and sparse SYK models for $6\leq N\leq 12$ Majorana fermions across a range of temperatures. Extending the framework developed in Ref. [28], this work investigates quantum hardware implementations for probing real-time evolution in the SYK model. However, this approach faces significant challenges, including the rapid growth in the number of parameters and controlled-NOT (CNOT) gates with increasing $N$. These limitations restricted simulations to $N\leq 12$, as larger systems incurred prohibitive computational overhead.
Figure 2: The CNOT gates in the PQC2 are $10^{2}$-fold improved beyond $N=12$ if the thermal state is prepared in the reinforcement learning framework compared to the real-time dynamics obtained by a first-order Trotterization [28]. The resource count is obtained in a noiseless scenario where both the environment and the agent are noiseless. Using the reward signal presented in Eq. 4 (where the signal constitutes the error in free energy, hence “Free energy reward”) and in Eq. 18 (where the signal constitutes the error in free energy and the thermal state fidelity, hence the “Free energy and fidelity reward”) and for different temperatures ($1/\beta$). The “ours” denotes the variational thermal state preparation under the reinforcement learning framework in Fig. 1. In Appendix A we further analyze the scaling of CNOT with the number of qubits and show that for our method the CNOT scales polynomially.
In this work, we address these challenges by integrating reinforcement learning (RL) with convolutional neural networks (CNNs) [29] into the variational quantum algorithm (VQA) framework. RL has recently emerged as a powerful tool for addressing quantum computing problems on near-term quantum devices [30, 31, 32, 33, 34, 35, 36]. In this approach, quantum circuits are modelled as sequences of actions governed by a trainable policy. The VQA cost function, independently optimized using a classical optimizer, provides periodic feedback that contributes to the final reward signal. This reward signal updates the policy, maximizing expected returns and guiding the selection of optimal actions for subsequent steps. For a brief review of RL for quantum circuit optimization check ref. [37].
@@
The remainder of the sections is organized as follows: in Sec. 2 we elaborate on the preliminaries, such as the problem statement and the variational thermal state preparation algorithm. This is followed by an introduction to the variational preparation of the thermal states in the RL framework in Sec. 3. In Sec. 4 we thoroughly discuss the main results of the paper, including the efficiency of the RL framework in noiseless and realistic noisy scenarios. Finally, in Sec. 5 we conclude the paper and discuss possible future work following the success of this paper.
@@ 2 Preliminaries
The dense Sachdev-Ye-Kitaev (SYK) model describes $N$ Majorana fermions with all-to-all $q$-body interactions governed by the Hamiltonian:
$H=\frac{(i)^{q/2}}{q!}\sum_{1\leq i_{1}<i_{2}<\cdots<i_{q}\leq N}J_{i_{1}i_{2}\ldots i_{q}}\chi_{i_{1}}\chi_{i_{2}}\cdots\chi_{i_{q}},$ (1)
where the Majorana operators $\chi_{i}$ satisfy $\{\chi_{i},\chi_{j}\}=\delta_{ij}$, and the antisymmetric random couplings $J_{i_{1}\ldots i_{q}}$ are drawn from a Gaussian distribution possessing zero mean and variance $\langle J_{i_{1}\ldots i_{q}}^{2}\rangle\propto J^{2}/N^{q-1}$.
$\overline{J_{i_{1}i_{2}i_{3}i_{4}}}=0,\quad\overline{J_{i_{1}i_{2}i_{3}i_{4}}^{2}}=\frac{3!}{N^{3}}\quad(J=1).$ (2)
In this work, we focus on the case $q=4$, where the interactions are quartic and randomly distributed. The dimension of the Hilbert space is $\dim(\mathcal{H})=2^{N/2}$, where $n=N/2$ is the number of qubits. This follows from the fact that two Majorana fermions can be represented by one complex spinless fermion, which corresponds to a single qubit. While the SYK model is analytically solvable in the large-$q$ limit [10], we constrain this research to $q=4$ as provided in Eq. (1). For simplicity, we denote $\beta J$ as $\beta$. The SYK model’s dense interaction structure presents formidable computational barriers rooted in its system-size scaling behavior. For quartic couplings ($q=4$), the Hamiltonian term count scales combinatorially as $\mathcal{O}(N^{4})$, driving exponentially growing resource requirements for classical simulations. This scaling collapse motivates the pursuit of quantum advantage in simulating such correlated fermionic systems. This scaling leads to two primary challenges (1) in classical simulation: The storage requirements grow as $\binom{N}{4}$, making exact diagonalization infeasible for large $N$. Meanwhile, in Quantum simulations, the number of gates required for parameterized quantum circuits (PQCs) increases polynomially with $N$, with a rapid growth in CNOT gate count. To address these challenges, sparsified versions of the SYK model have been proposed in prior works [40, 41], which reduce the number of terms in the Hamiltonian while approximating its essential physics. However, even with sparsification, quantum simulations remain resource-intensive. For instance, prior work on thermal state preparation for the SYK model using variational quantum algorithms was limited to $N=12$ Majorana fermions due to these overheads [27]. In this work, we tackle these computational bottlenecks by employing a reinforcement learning (RL) framework with an $\epsilon$-greedy policy. Our RL agent is implemented using a 3D convolutional neural network (CNN) that encodes quantum circuits as input tensors. This approach dynamically explores and optimizes circuit architectures while reducing CNOT gate overhead by over two orders of magnitude compared to standard PQCs arising due to the real-time dynamics of SYK models. Consequently, our method enables simulations of systems with more than $N=12$ Majorana fermions while preserving model fidelity.
@@ 2.1 Variational thermal state preparation
Thermal density matrices $\rho_{\beta}\propto e^{-\beta H}$ play a fundamental role in characterizing equilibrium properties of quantum systems at finite temperatures. Their preparation poses greater complexity compared to limiting cases: ground states ($\beta\to\infty$, pure states) and infinite-temperature states ($\beta\to 0$, maximally mixed), which admit efficient classical representations. The intermediate regime of finite $\beta$ proves particularly crucial for studying strongly correlated systems like the Sachdev-Ye-Kitaev model, where thermalization dynamics and holographic signatures emerge through entanglement structure and out-of-time-order correlations. To prepare $\rho_{\beta}$, we minimize the Helmholtz free energy $F$, given by $F=\langle H\rangle_{\beta}-TS_{v}$, with the temperature is given by $1/\beta$, $\langle H\rangle_{\beta}=\text{Tr}[\rho_{\beta}H]$ is the energy expectation value, where $\rho_{\beta}$ is the thermal state at temperature $T$, and ${S}_{v}=-\text{Tr}[\rho_{\beta}\log\rho_{\beta}]$ is the von Neumann entropy. We adopt the algorithm from Ref. [42] whose subroutines are defined in the following. Appendix C contains a comprehensive discussion of the theoretical foundations that explain why the variational thermal state preparation algorithm is effective.
@@
  • Measurement and entropy computation: First, implement the $\mathrm{PQC}_{1}(\boldsymbol{\theta})$ with tunable parameters $\boldsymbol{\theta}\in\mathbb{R}^{d}$. This prepares the variational state:
    $\lvert\psi(\boldsymbol{\theta})\rangle=\mathrm{PQC}_{1}(\boldsymbol{\theta})\lvert 0\rangle^{\otimes n},$
    where $n$ denotes the number of qubits. Subsequent computational basis measurements yield probability distributions $\{p_{i}(\boldsymbol{\theta})\}_{i=1}^{2^{n}}$, where $p_{i}(\boldsymbol{\theta})=|\langle i\,|\,\mathopen{}\psi(\boldsymbol{\theta})\rangle|^{2}$ corresponds to the measurement outcome probability for basis state $\lvert i\rangle$. The system’s Shannon entropy is then computed as:
    $S(\boldsymbol{\theta})=-\sum_{i=1}^{2^{n}}p_{i}(\boldsymbol{\theta})\ln p_{i}(\boldsymbol{\theta}),$
    quantifying the information entropy of the prepared quantum state. An elaborated discussion of the resource requirements in entropy estimation is provided in Appendix. B
@@
  • Initialization: After measuring PQC1 in the computational basis, the state reduces to:
    $\rho_{PQC1}=\sum_{i}p_{i}(\vec{\theta})|b_{i}\rangle\langle b_{i}|,$ (3)
    which works as the initial state for PQC2.
  • Ansatz construction: Employ a problem-specific or quantum hardware-inspired ansatz [43] with gates either derived from first-order Trotterization or obtained from the native gateset from quantum hardware.
@@
    $\rho_{\mathrm{PQC}_{2}}=\sum_{i}|a_{i}({\vec{\theta}})|^{2}\lvert\Psi_{i}({\vec{\phi}})\rangle\langle\Psi_{i}({\vec{\phi}})\rvert,$
    where $\lvert\Psi_{i}({\vec{\phi}})\rangle=\text{PQC}_{2}({\vec{\phi}})\lvert b_{i}\rangle$.
@@ 3 Variational thermal state preparation with reinforcement learning
We give an overview of the variational thermal state preparation of the SYK model within the reinforcement learning (RL) framework, for details of the framework see Appendix D. The algorithm is illustrated in Fig. 1. The framework is used to estimate the parameters of the PQC1 and the structure as well as parameters of the PQC2 elaborated in Sec. 2.1. Wherein we present state, the action representations, and the reward function used in this work.
@@
In the hybrid quantum-classical environment, the agent begins each episode with PQC1 (which includes a layer of RZ-RY-RZ rotations on each qubit followed by CNOT entangling gate connected in cyclic manner), then it sequentially adds gates to the PQC2 (which starts as an empty circuit at the beginning of each episode) until the maximum number of actions is reached. At each step, the state is represented by the current circuit, i.e., PQC1 combined with PQC2, and the action is to append a gate. To enable deep RL methods, states and actions are encoded in a format suitable for neural networks. Specifically, each state is represented by a tensor-based binary encoding of the circuit’s gate structure, as described in Appendix E. The action space consists of CNOT gates and 1-qubit parameterized rotation gates (RX, RY, and RZ). Notably, we omit the continuous parameters of the rotation gates, focusing instead on the estimated energy of the circuit to define the state representation for the RL agent.
@@
The reinforcement learning (RL) agent utilizes a 3D convolutional neural network (CNN), as the quantum circuit is encoded as a 3D tensor (refer to Appendix E). In prior studies [31, 44, 30, 45], this 3D representation was typically flattened into a 1D format to be processed by a fully connected feedforward neural network (FNN) [38, 39]. However, flattening 3D data into 1D can result in the loss of critical spatial information, diminished model expressiveness, and potentially suboptimal performance. Architectures like 3D CNNs that can directly process the 3D structure are often preferable when working with volumetric data. If not flattened when operating on a structure like a 3D-CNN [46, 47], the neural net receives a dataset with distinguishable spatial information. In the case of the 3D binary encoding, where the X-axis represents the position of gates in PQC2, the Y-axis denotes the type of gate, and the Z-direction is the depth of PQC2. In Appendix. F we elaborately show that directly feeding higher-dimensional data through a 3D-CNN provides improved training, leading to lower approximation error with more compact parameterized quantum circuits compared to processing flattened data through an FNN. The hypermeters settings of the agent is discussed in Appendix I.
@@
$r=\begin{cases}5,&\text{if }F_{t}(\beta=\text{const.},\vec{\theta},\vec{\phi})\leq\zeta_{F},\\ -5,&\text{if }F_{t}(\beta=\text{const.},\vec{\theta},\vec{\phi})>\zeta_{F}\;\text{and,}\;\text{step no.}=D_{\text{max}},\\ E_{\text{term}},&\text{otherwise}.\end{cases}$ (4)
@@
In Sec. 4 we show that using the reward in Eq. 4 we can achieve a very good approximation of the free energy, but it cannot provide as good an approximation to the energy expectation of the SYK Hamiltonian and the entropy. Hence, we engineer the reward signal to incorporate the accuracy of the prepared thermal state by calculating its fidelity from the target thermal state after PQC2. Hence, the modified reward function is given by
$r=\begin{cases}5,&\text{if }F_{t}(\beta=\text{const.},\vec{\theta},\vec{\phi}){\leq}\zeta_{F}\;\text{and,}\;\textrm{Fid}(\rho({\vec{\phi}}))\geq\zeta_{\textrm{Fid.}}\\ -5,&\text{if }\textrm{Fid}(\rho({\vec{\phi}}))<\zeta_{\textrm{Fid}}\;\text{and,}\;\text{step no.}=D_{\text{max}},\\ 0.6\times E_{\text{term}}+0.4\times\textrm{Fid}_{\text{term}},&\text{otherwise},\end{cases}$ (5)
Where $\textrm{Fid}_{\text{term}}=2\times\textrm{Fid}(\rho({\vec{\phi}}))-1$, $\textrm{Fid}(\rho({\vec{\phi}}))={\mbox{Tr}}\Big{(}\sqrt{\rho_{\beta}^{1/2}\rho(\vec{\phi})~{}\rho_{\beta}^{1/2}}\Big{)}$ and $\rho_{\beta}$ represents the density matrix that corresponds to the true thermal state. $\zeta_{\textrm{Fid}}$ is the user-defined threshold on fidelity, throughout the paper, we choose $\zeta_{\textrm{Fid}}=0.9$. An elaborated analysis of the reward function defined in Eq. 18 is provided in the Appendix J.